# Masked Generative Extractor for Synergistic Representation and 3D Generation of Point Clouds

### v1

@@ Abstract
In the field of 2D image generation modeling and representation learning, Masked Generative Encoder (MAGE) has demonstrated the synergistic potential between generative modeling and representation learning. Inspired by this, we propose Point-MAGE to extend this concept to point cloud data. Specifically, this framework first utilizes a Vector Quantized Variational Autoencoder (VQVAE) to reconstruct a neural field representation of 3D shapes, thereby learning discrete semantic features of point patches. Subsequently, by combining the masking model with variable masking ratios, we achieve synchronous training for both generation and representation learning. Furthermore, our framework seamlessly integrates with existing point cloud self-supervised learning (SSL) models, thereby enhancing their performance. We extensively evaluate the representation learning and generation capabilities of Point-MAGE. In shape classification tasks, Point-MAGE achieved an accuracy of 94.2% on the ModelNet40 dataset and 92.9% (+1.3%) on the ScanObjectNN dataset. Additionally, it achieved new state-of-the-art performance in few-shot learning and part segmentation tasks. Experimental results also confirmed that Point-MAGE can generate detailed and high-quality 3D shapes in both unconditional and conditional settings.
@@ 1 Introduction
Recent studies have confirmed that representation learning frameworks based on masked modeling Yu_2022_CVPR ; liu2022masked ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 can extract high-quality 3D representations. At the same time, generative models zhang20223dilg ; Mittal_2022_CVPR ; Yan_2022_CVPR ; zhang20233dshape2vecset have also made significant achievements in generating realistic 3D shapes. However, there is little research that attempts to combine the potential of these two types of models to explore their joint application in 3D shapes. We note that in the field of 2D images, Masked Generative Encoder (MAGE) Li_2023_CVPR proposed an innovative unified framework for generation training and representation learning, and experimentally verified the synergistic effects between these two tasks. This inspired us to pose a question: Can the characteristics demonstrated in the MAGE framework be extended to point cloud data to achieve deeper 3D representation learning and generation capabilities?
The strategy of MAGE Li_2023_CVPR is to combine Masked Image Modeling (MIM) with variable masking ratios. Specifically, this method conducts generation training under high masking ratios, while representation learning is performed under low masking ratios. Following this idea, we explore combining masked point cloud modeling with variable masking ratios to realize the Point-MAGE framework. However, due to the fundamental differences in structure and distribution between images and point cloud data, this direct transfer faces several challenges.
@@
Firstly, MAGE Li_2023_CVPR effectively obtains discrete semantic features from images through pixel-wise reconstruction training. However, directly using point clouds as reconstruction targets introduces sampling bias. This is because 3D shapes are inherently continuous, while point clouds obtained through sampling merely represent them discretely. This leads to non-zero loss between different point cloud samples of the same object, affecting the accurate understanding of shapes. Moreover, methods solely relying on point cloud reconstruction fail to acquire more advanced representations of 3D shapes, such as neural radiance fields or triangle meshes, thereby limiting the potential of models in generation tasks. Therefore, simply applying variable masking ratios to existing point cloud SSL methods Yu_2022_CVPR ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 cannot achieve the goals of Point-MAGE, as these existing methods typically use Chamfer distance (CD) as the reconstruction loss for point cloud reconstruction. To address these challenges, we instead focus the reconstruction targets on 3D shapes represented as neural radiance fields (NeRF) and extract discrete semantic features of point clouds through a VQVAE van2017neural . Point-MAGE leverages these high-level semantic discrete features as input and reconstruction targets, learning the probability distribution of masked tokens. After training, our method not only iteratively performs three-dimensional shape generation tasks but also resolves sampling bias in representation learning through the vector quantization process. This is because different point cloud samples from the same local block are unified into the same discrete token during the vector quantization stage, ensuring consistency of representations.
@@
Secondly, the inherent unordered nature of point cloud poses a challenge for Transformer-based generative models, as these models require serialized inputs. To address this issue, we partition the complete point cloud into multiple patches, each defined by a fixed number of points and a central point. To serialize these patches, we adopt the Morton sorting algorithm used in PointGPT NEURIPS2023_5ed5c3c8 , which can sort point cloud patches while preserving geometric relationships. In generation task, since the central point of point cloud patch is unknown, we introduce relative position encoding in the generator of Point-MAGE as a replacement for traditional absolute position encoding, and use the central points as the reconstruction target.
Finally, we conducted an analysis of the synergistic effects of the two tasks. We found that generative models benefit from the variable masking ratio. Unlike traditional methods that reconstruct all tokens from scratch, our approach leverages a small number of visible tokens to provide additional sample information, which not only accelerates convergence but also enhances the model’s final performance. For representation learning, by shifting the reconstruction target from raw point clouds to discrete semantic features, our framework successfully overcomes the variance issue in point cloud sampling and focuses more on the global features of 3D shapes. Furthermore, the Point-MAGE framework can be integrated with SSL model designs proposed in previous work, further enhancing the performance of representation learning. Experimental results demonstrate the effectiveness of this strategy, with Point-MAGE achieving state-of-the-art (SOTA) performance on multiple downstream tasks. Specifically, our method achieved an accuracy of 94.2% on the ModelNet40 dataset Wu_2015_CVPR and 92.9% on the ScanObjectNN dataset yi2016scalable , respectively, for shape classification tasks. In the part segmentation task on the ShapeNetPart yi2016scalable dataset, our method achieved a mean class Intersection over Union (IoU) of 85.0%. Additionally, we demonstrated the pre-trained model’s ability to generate realistic 3D shapes in both unconditional and conditional settings, further showcasing the flexibility and powerful generative capabilities of Point-MAGE.
@@
Our main contributions can be summarized as follows: (I) We propose Point-MAGE, a unified framework for point cloud data representation learning and generative training that combines a two-stage training method with a variable masking ratio design. (II) Point-MAGE overcomes the variance issue in point cloud sampling and can be integrated with model architectures from previous SSL works, setting a new SOTA benchmark for downstream tasks such as shape classification, few-shot learning, and part segmentation. (III) Point-MAGE is capable of generating high-quality 3D shapes in both unconditional and conditional settings.
@@ 2 Related Works
In recent years, self-supervised frameworks for point cloud representation learning have made significant progress, forming two main model paradigms. Firstly, methods based on contrastive learning Yang_2018_CVPR ; xie2020pointcontrast ; Navaneet_2020_CVPR ; Zhang_2021_ICCV ; Afham_2022_CVPR have played an important role in point cloud self-supervised learning (SSL). These methods can explore the intrinsic properties of point clouds, but they often rely on cross-modal information or teacher models, which undoubtedly add extra burden to the training process. Secondly, more closely related to our work is another class of methods based on masked point cloud modeling. Inspired by the successful experience of masked modeling in natural language processing (NLP) radford2018improving ; devlin2018bert ; radford2019language ; brown2020language and image processing he2022masked ; Wei_2022_CVPR ; zhou2021ibot fields, recent research has focused on randomly masking point cloud regions and reconstructing these regions Yu_2022_CVPR ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 , prompting the encoder to learn high-quality point cloud representations. In particular, Point-M2AE zhang2022point proposed an innovative hierarchical transformer structure, extending Point-MAE pang2022masked to capture multi-scale features of point clouds. On the other hand, PointGPT attempts to solve the problem of global shape leakage through autoregressive generation, but it does not have the ability to generate high-quality 3D models from scratch. Additionally, methods based on masked region reconstruction are susceptible to point cloud sampling variance, which may introduce additional understanding and bias issues. To overcome these challenges, our method adopts a strategy of masked modeling in the high-level semantic space, effectively addressing the aforementioned issues.
@@
In the field of 3D generation, Generative Adversarial Networks (GANs) wu2016learning ; achlioptas2018learning ; Chen_2019_CVPR ; Ibing_2021_CVPR ; zheng2022sdf constitute a major generative paradigm, despite their limitations in training stability and mode collapse issues. Meanwhile, some research has turned to alternative approaches such as Normalizing Flows (NFs) dinh2014nice ; rezende2015variational ; yang2019pointflow , Energy-Based Models (EBMs) lecun2006tutorial ; xie2016theory ; Xie_2021_CVPR ; xie2020generative , and Autoregressive Models (ARs) Mittal_2022_CVPR ; cheng2022autoregressive ; Yan_2022_CVPR ; zhang20223dilg . Recently, inspired by the significant progress of Diffusion Models (DMs) dhariwal2021diffusion ; song2020score in image synthesis, some studies have begun to explore the application of diffusion models to 3D shape generation vahdat2022lion ; hui2022neural ; Shue_2023_CVPR ; Chou_2023_ICCV ; zhang20233dshape2vecset . Although these models can generate realistic 3D models, they often lack the ability to extract high-quality semantic representations from 3D shapes, limiting their direct application in point cloud representation learning tasks. To address this challenge, we propose an innovative approach aimed at generating diverse samples while extracting high-quality semantic embeddings from 3D shapes to promote data generalization and enhance the performance of downstream tasks. We draw on the successful experience of MAGE Li_2023_CVPR in image synthesis and representation learning, and realize the organic combination of generation and representation learning on point clouds through a two-stage scheme. In the first stage, we use VQVAE van2017neural to train an encoder and decoder to obtain latent codes. Subsequently, in the second stage, we further train the extractor for representation learning through masked modeling and use a bidirectional transformer generator to complete the generation task.
@@ 3 Method
Point-MAGE is designed as a unified framework for both point cloud generation tasks and representation learning. As shown in Figure 1, the framework first divides the input point cloud into multiple local point blocks and serializes them based on their geometric positional relationships. Then, it uses a mini-PointNet Qi_2017_CVPR model for initial feature extraction of each point block. Subsequently, by training a VQVAE van2017neural model, these point block features are quantized into discrete semantic latents. Finally, the framework combines a variable masking ratio with a Vision Transformer (ViT)-based extractor-generator architecture to predict masked semantic tokens.
Figure 1: Overall pipeline of Point-MAGE. The training process consists of two stages. In the first stage, a VQVAE van2017neural is used to reconstruct 3D shapes represented by NeRF, converting the input point cloud into a series of discrete semantic latents. The second stage is a masked token reconstruction process, where a ViT-based extractor-generator architecture is employed to extract high-quality feature representations from the unmasked tokens and to reconstruct the masked tokens.
@@ 3.1 Point Patch Embedding
To provide serialized input, we follow NEURIPS2023_5ed5c3c8 and encode the coordinates of the central points $x_{i}$ into one-dimensional space using Morton ordering morton1966computer , and then sort them to obtain serialized central points $C^{s}\in\mathbb{R}^{G\times 3}$ and point cloud patches $P^{s}\in\mathbb{R}^{G\times 3}$. We project each patch into an embedding vector using the mini-PointNetQi_2017_CVPR module,
[Uncaptioned image]${\color[rgb]{0.875,0.4765625,0.3671875}~{}:P^{s}_{E}}=\{e_{i}\}_{i\in\mathcal{G}}=\text{PointNet}\left(P^{s}\right),~{}P^{s}_{E}\in\mathbb{R}^{G\times D},$ (3)
@@ 3.2 Tokenization
We train a VQVAE van2017neural to obtain a quantizer that embeds input point patches into a series of semantic latents. This tokenization step enables our model to focus on high-level semantics, which helps to overcome the variance issue in point cloud sampling and allows for a more comprehensive understanding of 3D shapes.
@@
We use a ViT-based encoder to learn the local latents $T$. The input to the transformer includes the sorted point block embeddings $P^{s}_{E}$ and the sinusoidal positional encodings vaswani2017attention $C^{s}_{E}$ of the center point,
$T=\{t_{i}\}_{i\in\mathcal{G}}=\text{Encoder}\left(P^{s}_{E},C^{s}_{E}\right),~{}T\in\mathbb{R}^{G\times D}.$ (4)
@@
NeRF Reconstruction.
We utilize a ViT-based decoder along with a reconstruction head to reconstruct 3D shapes represented in NeRF. The decoder takes as input the quantized latents and the position encoding of the center points,
$T^{D}=\{t^{d}_{i}\}_{i\in\mathcal{G}}=\text{Decoder}\left(T^{Q},C^{s}_{E}\right),~{}T^{D}\in\mathbb{R}^{G\times D}.$ (6)
@@
$\mathcal{O}(x)=\text{Sigmoid}\left(\text{MLP}\left(x,t^{d}_{x}\right)\right),~{}\mathcal{O}(x)\in\left[0,1\right].$ (7)
@@
$\mathcal{L}_{n}=\mathbb{E}_{x\in\mathbb{R}^{3}}[\text{BCE}(\mathcal{O}(x),\hat{\mathcal{O}}(x))]+\lambda\mathbb{E}_{x\in\mathbb{R}^{3}}[\mathbb{E}_{i\in\mathcal{M}}\|\epsilon(t^{q}_{i})-t_{i}\|^{2}],$ (8)
where $\hat{\mathcal{O}}(x)$ denotes the ground truth, and $\epsilon(\cdot)$ represents the stop-gradient operation. In this paper, $\lambda=0.25$ is a weighting coefficient.
@@ 3.3 Pre-training
After obtaining the tokenizer, we simultaneously train an extractor and a generator model through masking modeling. Below, we introduce the masking strategy, model design, and optimization objectives of Point-MAGE.
@@
For a given point cloud $X\in\mathbb{R}^{n\times 3}$, we first use the trained tokenizer to obtain the corresponding discrete tokens $T^{Q}$. Subsequently, we introduce a variable mask ratio to apply masking, which can be calculated as:
$m_{r}\sim N(\mu=0.55,\sigma^{2})~{}\text{with}~{}m_{r}\in\left[0.5,1\right]$ (9)
Where $N$ represents the truncated Gaussian distribution. After masking, to reduce training time and memory consumption, we randomly drop a portion of masked tokens, ensuring that only $0.5\times G$ tokens are input to the extractor, following the MAGE Li_2023_CVPR .
@@
$T^{E}=\{t^{e}_{i}\}_{i\in\mathcal{G},m_{i}=0}=\text{Extractor}(T^{Q}_{m},C^{s}_{E}),~{}T^{E}\in\mathbb{R}^{\frac{G}{2}\times D},$ (10)
@@
$T^{G}=\{t^{g}_{i}\}_{i\in\mathcal{G}}=\text{Generator}(T^{E}_{p},C^{s}_{RE},\mathcal{C}),~{}T^{G}\in\mathbb{R}^{G\times D},$ (11)
@@
$\mathcal{L}_{rt}=-\mathbb{E}_{t^{q}_{i}\in\mathcal{Q}}[\mathop{\sum}\limits_{\forall i,m_{i}=1}\log p(t^{q}_{i}|T^{G}_{m})]+\frac{1}{|M|}\mathop{\sum}\limits_{\forall i,m_{i}=1}\|c_{i},\hat{c}_{i}\|^{2}$ (12)

### v2

@@ Abstract
Representation and generative learning, as reconstruction-based methods, have demonstrated their potential for mutual reinforcement across various domains. In the field of point cloud processing, although existing studies have adopted training strategies from generative models to enhance representational capabilities, these methods are limited by their inability to genuinely generate 3D shapes. To explore the benefits of deeply integrating 3D representation learning and generative learning, we propose an innovative framework called Point-MGE. Specifically, this framework first utilizes a vector quantized variational autoencoder to reconstruct a neural field representation of 3D shapes, thereby learning discrete semantic features of point patches. Subsequently, we design a sliding masking ratios to smooth the transition from representation learning to generative learning. Moreover, our method demonstrates strong generalization capability in learning high-capacity models, achieving new state-of-the-art performance across multiple downstream tasks. In shape classification, Point-MGE achieved an accuracy of 94.2% (+1.0%) on the ModelNet40 dataset and 92.9% (+5.5%) on the ScanObjectNN dataset. Experimental results also confirmed that Point-MGE can generate high-quality 3D shapes in both unconditional and conditional settings.
@@ 1 Introduction
Recent studies have confirmed that masked modeling-based representation learning frameworks Yu et al. (2022); Pang et al. (2022); Zhang et al. (2022b) can extract high-quality 3D representations. At the same time, generative modelss Zhang et al. (2022a); Mittal et al. (2022); Zhang et al. (2023a) have achieved remarkable success in generating realistic 3D shapes. Nevertheless, few studies have explored the deep integration of these two models to unlock their combined potential in 3D shape processing. In natural language processing (NLP) Touvron et al. (2023) and 2D image domains Li et al. (2023), we have witnessed the versatility of pre-trained models: not only can they be used for diverse generative tasks such as question answering (QA), code generation, and image generation, but they can also serve as feature extractors for various downstream tasks through fine-tuning. These achievements prompt us to consider: can the benefits of this model integration be similarly applied to the representation of point cloud data and 3D shape generation? To this end, we conducted an in-depth analysis of the characteristics of both models and found that masked reconstruction methods offer a potential pathway to achieve this integration, though several challenges remain.
@@
Firstly, generative learning is easily trained on discrete distributions, whereas existing point cloud representation learning Pang et al. (2022); Zhang et al. (2022b) tends to operate within continuous feature spaces. In prior work, Point-Bert Yu et al. (2022) learns a tokenizer by reconstructing the input point cloud and optimizing the chamfer distance (CD) loss. However, using point clouds directly as the reconstruction target introduces a sampling bias. This is because 3D shapes are inherently continuous, but point clouds sampled from these shapes can only approximate them in a discrete form. This sampling bias can lead to non-zero loss differences between different point cloud samples of the same object, thereby affecting the precise understanding of shape features. Moreover, methods that rely solely on point cloud reconstruction are limited in capturing high-level representations of 3D shapes. For instance, they cannot generate richer shape descriptions such as neural radiance fields (NeRFs) or triangular meshes, which restricts the model’s potential in generative tasks. To address this, we shift the reconstruction target to 3D shapes represented by NeRFs and utilize a vector quantized variational autoencoder (VQVAE) Van Den Oord et al. (2017) to extract discrete semantic features from point clouds. The Point-MGE framework leverages these high-level semantic discrete features not only as input but also as reconstruction targets, thereby learning the probability distribution of masked tokens. During the vector quantization process, different point cloud samples from the same local region are consistently mapped to the same discrete tokens, ensuring consistency and robustness in the representations.
@@
Secondly, the inherent unordered nature of point cloud poses a challenge for Transformer-based generative models, as these models require serialized inputs. To address this issue, we partition the complete point cloud into multiple patches, each defined by a fixed number of points and a central point. To serialize these patches, we adopt the Morton sorting algorithm used in PointGPT Chen et al. (2023), which can sort point cloud patches while preserving geometric relationships. In generation task, since the central point of point cloud patch is unknown, we introduce relative position encoding in the generator of Point-MGE as a replacement for traditional absolute position encoding, and use the central points as the reconstruction target.
Finally, generative training tends to adopt higher masking ratios, or even full masking, whereas representation learning typically uses relatively low masking ratios. In 2D image processing, MAGE Li et al. (2023) achieves variable masking ratios by randomly sampling them from a Gaussian distribution, utilizing these variable ratios to perform the task. However, due to the irregularity of point cloud data and the lack of support from large-scale datasets, such drastic fluctuations in masking ratios may adversely affect model convergence. To address this issue, we have carefully designed a sliding masking ratio mechanism to smoothly adjust the masking ratio during training. Specifically, the masking ratio gradually increases as training iterations progress, facilitating a smooth transition from representation learning to generative training. We also conducted a comparative analysis of the effects of different sliding curves. Point-MGE can scale to high-capacity model training and has demonstrated outstanding performance across multiple benchmark datasets. It achieved an accuracy of 94.2% on the ModelNet40 dataset Wu et al. (2015) for shape classification and 92.9% on the ScanObjectNN dataset Uy et al. (2019). Additionally, in the part segmentation task on the ShapeNetPart dataset Yi et al. (2016), our method attained an average class Intersection over Union (IoU) of 85.0%. Furthermore, we also showcased the capability of the pre-trained model to generate realistic 3D shapes in both unconditional and conditional settings.
@@
Our main contributions can be summarized as follows: (I) We propose Point-MGE, a unified framework for point cloud data representation learning and generative training that combines a two-stage training method with a sliding masking ratio design. (II) Point-MGE overcomes the variance issue in point cloud sampling and scales to high-capacity model training, setting a new SOTA benchmark for downstream tasks. (III) Point-MGE is capable of generating high-quality 3D shapes in both unconditional and conditional settings.
@@ 2 Related Works
In recent years, self-supervised frameworks for point cloud representation learning have made significant progress, forming two main model paradigms. Firstly, methods based on contrastive learning Xie et al. (2020b); Zhang et al. (2021); Afham et al. (2022) have played an important role in point cloud self-supervised learning. These methods can explore the intrinsic properties of point clouds, but they often rely on cross-modal information or teacher models, which undoubtedly add extra burden to the training process. Secondly, inspired by natural language processing Devlin et al. (2018); Brown et al. (2020) and 2D computer vision He et al. (2022); Wei et al. (2022), a more closely related area to our work is another category of methods based on masked point cloud modeling Yu et al. (2022); Pang et al. (2022); Zhang et al. (2022b); Chen et al. (2023). Point-M2AE Zhang et al. (2022b) proposed an innovative hierarchical transformer structure, extending Point-MAE Pang et al. (2022) to capture multi-scale features of point clouds. On the other hand, PointGPT Chen et al. (2023) attempts to solve the problem of global shape leakage through autoregressive generation, but it does not have the ability to generate high-quality 3D models from scratch. Additionally, methods based on masked region reconstruction are susceptible to point cloud sampling variance, which may introduce additional understanding and bias issues. To overcome these challenges, our method adopts a strategy of masked modeling in the high-level semantic space, effectively addressing the aforementioned issues.
@@
In the field of 3D generation, generative adversarial networks (GANs) Achlioptas et al. (2018); Chen and Zhang (2019); Ibing et al. (2021); Zheng et al. (2022) constitute a major generative paradigm, despite their limitations in training stability and mode collapse issues. Meanwhile, some research has turned to alternative approaches such as normalizing flows (NFs) Dinh et al. (2014); Rezende and Mohamed (2015); Yang et al. (2019), energy-based models (EBMs) Xie et al. (2016, 2020a, 2021), and autoregressive models (ARs) Mittal et al. (2022); Cheng et al. (2022); Yan et al. (2022); Zhang et al. (2022a). Recently, inspired by the significant progress of diffusion models (DMs) Dhariwal and Nichol (2021) in image synthesis, some studies have begun to explore the application of diffusion models to 3D shape generation Zhou et al. (2021); Vahdat et al. (2022); Hui et al. (2022); Shue et al. (2023); Chou et al. (2023); Zhang et al. (2023a) . Although these models can generate realistic 3D models, they often lack the ability to extract high-quality semantic representations from 3D shapes, limiting their direct application in point cloud representation learning tasks. To address this challenge, we propose an innovative approach aimed at generating diverse samples while extracting high-quality semantic embeddings from 3D shapes to promote data generalization and enhance the performance of downstream tasks.
@@ 3 Method
Point-MGE is designed as a unified framework for both 3D generation tasks and representation learning. As shown in Figure 1, the framework first divides the input point cloud into multiple local point patches and serializes them based on their geometric positional relationships. Then, it uses a mini-PointNet Qi et al. (2017a) model for initial feature extraction of each point patch. Subsequently, by training a VQVAE model, these features are quantized into discrete tokens. Finally, the framework combines a variable masking ratio with a vision transformer (ViT)-based extractor-generator architecture to predict masked tokens.
Figure 1: Overall pipeline of Point-MGE. First, a VQVAE is used to reconstruct 3D shapes represented by NeRFs, converting the input point cloud into a series of discrete semantic tokens. Then, where a ViT-based extractor-generator architecture is employed to extract high-quality feature representations from the unmasked tokens and to reconstruct the masked tokens.
@@ 3.1 Point Patch Embedding
To provide serialized input, we follow Chen et al. (2023) and encode the coordinates of the central points $x_{i}$ into one-dimensional space using morton ordering Morton (1966), and then sort them to obtain serialized central points $C^{s}\in\mathbb{R}^{G\times 3}$ and point cloud patches $P^{s}\in\mathbb{R}^{G\times 3}$. We project each patch into an embedding vector using the mini-PointNetQi et al. (2017a) module,
$P^{s}_{E}=\{e_{i}\}_{i\in\mathcal{G}}=\text{PointNet}\left(P^{s}\right),~{}P^{s}_{E}\in\mathbb{R}^{G\times D},$ (3)
@@ 3.2 Tokenization
We train a VQVAE Van Den Oord et al. (2017) to obtain a quantizer that embeds input point patches into a series of semantic tokens. This tokenization step enables our model to focus on high-level semantics, which helps to overcome the variance issue in point cloud sampling and allows for a more comprehensive understanding of 3D shapes.
@@
We use a ViT-based encoder to learn the local latents $T$. The input to the transformer includes the sorted point block embeddings $P^{s}_{E}$ and the sinusoidal positional encodings Vaswani et al. (2017) $C^{s}_{E}$ of the center point,
$T=\{t_{i}\}_{i\in\mathcal{G}}=\text{Enc}\left(P^{s}_{E},C^{s}_{E}\right),~{}T\in\mathbb{R}^{G\times D}.$ (4)
@@
Reconstruction.
We utilize a ViT-based decoder to reconstruct the 3D shape’s NeRFs $\hat{\mathcal{O}}(x)\in\{0,1\}$, where $\hat{\mathcal{O}}(x)=1$ indicates that the point $x$ in space belongs to the object. The decoder takes as input the quantized tokens and the position encoding of the center points,
$T^{D}=\{t^{d}_{i}\}_{i\in\mathcal{G}}=\text{Dec}\left(T^{Q},C^{s}_{E}\right),~{}T^{D}\in\mathbb{R}^{G\times D}.$ (6)
@@
$\mathcal{O}(x)=\text{Sigmoid}\left(\text{MLP}\left(x,t^{d}_{x}\right)\right).$ (7)
@@
$\footnotesize\begin{split}\mathcal{L}_{n}=\mathbb{E}_{x\in\mathbb{R}^{3}}[\text{BCE}(\mathcal{O}(x),\hat{\mathcal{O}}(x))]+~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\\ \lambda\mathbb{E}_{x\in\mathbb{R}^{3}}[\mathbb{E}_{i\in\mathcal{M}}\|\epsilon(t^{q}_{i})-t_{i}\|^{2}],\end{split}$ (8)
where $\epsilon(\cdot)$ represents the stop-gradient operation. In this paper, $\lambda=0.25$ is a weighting coefficient.
@@ 3.3 Pre-training
After obtaining the tokenizer, we simultaneously train an extractor and a generator model through masking modeling. Below, we introduce the masking strategy, model design, and optimization objectives of Point-MGE.
Methods Reference Params(M) ModelNet40 ScanObjectNN
w/o Vote w/ Vote OBJ_BG OBJ_ONLY PB_T50_RS
Supervised Learning Only
PointNet Qi et al. (2017a) CVPR 2017 3.5 - 89.2 73.3 79.2 68.0
PointNet++ Qi et al. (2017b) Neurips 2017 1.5 - 90.7 82.3 84.3 77.9
DGCNN Wang et al. (2019) TOG 2019 1.8 - 92.9 82.8 86.2 78.1
PointMLP Ma et al. (2022) ICLR2022 12.6 94.1 94.5 - - 85.2
P2P-HorNet Wang et al. (2022) NeurIPS 2022 195.8 94.0 - - - 89.3
with Single-Modal Pre-Training
Point-BERT Yu et al. (2022) CVPR 2022 22.1 92.7 93.2 87.4 88.1 83.1
Mask-Point Liu et al. (2022) ECCV 2022 22.1 - 93.8 89.3 88.1 84.3
Point-MAE Pang et al. (2022) ECCV 2022 22.1 93.2 93.8 90.0 88.2 85.2
Point-M2AE Zhang et al. (2022b) Neurips 2022 15.3 93.4 94.0 91.2 88.8 86.4
PointGPT Chen et al. (2023) Neurips 2023 19.5 - 94.0 91.6 90.0 86.9
GPM Li et al. (2024) CVPR2024 41.5 94.0 94.3 90.2 90.0 84.8
Point-MGE - 15.3 93.8 94.2 92.9 91.0 87.1
Improvement (baseline: Point-BERT) - - +1.1 +1.0 +5.5 +2.9 +4.0
Point-MAE-B Pang et al. (2022) ECCV 2022 120.1 - 94.2 94.2 93.9 90.2
Point-M2AE-B Zhang et al. (2022b) Neurips 2022 77.5 - 94.3 95.2 94.3 91.2
PointGPT-B Chen et al. (2023) Neurips 2023 82.1 - 94.4 95.8 95.2 91.9
Point-MGE-B - 77.5 - 94.5 96.1 95.7 92.3
with Cross-Modal Pre-Training
ACT Dong et al. (2023) ICLR 2023 22.1 93.2 93.7 93.3 91.9 88.2
Joint-MAE Guo et al. (2023) IJCAI 2023 - - 94.0 90.9 88.9 86.1
I2P-MAE Zhang et al. (2023b) CVPR 2023 15.3 93.7 94.1 94.2 91.6 90.1
Recon Qi et al. (2023) ICML 2023 44.3 94.1 94.5 95.2 93.3 90.6
Table 1: Shape classification results on ModelNet40 Wu et al. (2015) and ScanObjectNN Uy et al. (2019). We report the accuracy (%) of both training from scratch and fine-tuning models on ModelNet40 and ScanObjectNN datasets.
@@
For a given point cloud $X\in\mathbb{R}^{n\times 3}$, we first use the trained tokenizer to obtain the corresponding discrete tokens $T^{Q}$. Subsequently, we introduce a sliding masking ratio to apply masking, which can be calculated as:
$m_{r}=\beta+(1-\beta)(\frac{\gamma}{\Gamma})^{u},~{}m_{r}\in\left[\beta,1\right],$ (9)
where, $\gamma$ and $\Gamma$ denote the current training iteration and the total number, respectively. The parameter $u>0$ is a hyperparameter used to control the trend of the sliding function. After masking, to reduce training time and memory consumption, we randomly drop a portion of masked tokens, ensuring that only $V=(1-\beta)\times G$ tokens are input to the extractor.
@@
$\footnotesize T^{E}=\{t^{e}_{i}\}_{i\in\mathcal{G},m_{i}=0}=\text{Ext}(T^{Q}_{m},C^{s}_{E}),~{}T^{E}\in\mathbb{R}^{V\times D},$ (10)
@@
$T^{G}=\{t^{g}_{i}\}_{i\in\mathcal{G}}=\text{Gen}(T^{E}_{p},C^{s}_{RE},\mathcal{C}),~{}T^{G}\in\mathbb{R}^{G\times D},$ (11)
@@
$\footnotesize\mathcal{L}_{rt}=-\mathbb{E}_{t^{q}_{i}\in\mathcal{Q}}[\mathop{\sum}\limits_{\forall i,m_{i}=1}\log p(t^{q}_{i}|T^{G}_{m})]+\frac{1}{|M|}\mathop{\sum}\limits_{\forall i,m_{i}=1}\|c_{i},\hat{c}_{i}\|^{2},$ (12)